# Practical multi-fidelity machine learning: fusion of deterministic and Bayesian models

### v1

@@ 3.1 Model for data-scarce or low-dimensional scenarios: KRR-LR-GPR
We argue that a logical LF model to pair with GPRs in a data-scarce MF scenario is kernel ridge regression (KRR), i.e. the deterministic formulation of GPR [16]. KRR is a kernel machine learning method that is a point estimate of a GPR with the same kernel, therefore also equivalent to a DNN with an infinitely wide hidden layer [55]. KRR is robust to noisy data, and easy to train due to having few hyperparameters. More importantly, existing KRR approaches can lower the time complexity to a range between $\mathcal{O}(n^{2})$ and $\mathcal{O}(n)$ by combining proper matrix decomposition technique for matrix inversion, thus handling larger datasets than Polynomial chaos expansion (PCE) or GPR [56, 57] – invaluable for LF datasets because they are typically larger than HF ones. Other LF models such as linear regression or deep neural networks (e.g., see Table 1) have important drawbacks. Linear regression models are sensitive to the choice of the basis functions (often considered to be polynomials whose order is a hyperparameter); this is why we argue for their use as a simple transfer-learning model $g(\mathbf{x})$ but not as a LF or residual model. Conversely, DNNs have a large number of hyperparameters and are therefore less practical to train. Good practice involves starting with simpler models such as KRR that are easier to train and only considering DNNs when the simpler methods fail. We note that GPRs can also be a good choice for a LF model if the LF dataset is small (typically below a few thousand points for most common hardware), but training and inference time increases when compared to KRR, as demonstrated later in this article.
@@
The proposed KRR-LR-GPR established on Eq. (6) is easy to implement and the main steps are listed in Algorithm 1, while additional details are provided in Appendix B. Recall that the key idea is to train a GPR model where, instead of considering a zero mean function, we assume it to be a linear model whose features are the LF model. This is advantageous as there is an explicit solution for finding the parameters $\boldsymbol{\rho}$ from the HF data, no longer treating them as hyperparameters: Step 2.1 in Algorithm 1 or Eq. (24) in the Appendix. The mean model of the GPR then captures most of the HF response by linear transfer-learning of the LF model, leaving the remaining nonparameteric approximation for the residual via a Gaussian process whose kernel hyperparameters are optimized as usual (Step 2.2 in Algorithm 1). The resulting KRR-LR-GPR model predicts the response and corresponding uncertainty, usually by following a Normal distribution $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{s}_{f^{h}}^{2}(\mathbf{x}))$ to ensure fast training and inference (exact integration of posterior and posterior predictive distributions). We find that a first-order linear transfer-learning model including a bias term, i.e. $\mathbf{m}(\mathbf{x})=[1,f^{l}(\mathbf{x})]^{T}$, provides robust results according to the ablation study for different problems (see Appendix E.1).
@@
Result: $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{s}_{f^{h}}^{2}(\mathbf{x}))$
@@ 3.2 Model for data-rich or high-dimensional scenarios: DNN-LR-BNN
Pairing a DNN as the LF model with a BNN in a MF model leverages the scalability of DNNs with the predictive abilities of BNNs. We propose two MF models of this type, and argue in favor of the simpler one. The simpler model summarized in Fig.2 is labeled DNN-LR-BNN, where transfer-learning is done via linear regression (LR) as in Eq. (4), and the residual is modeled by a BNN. This can be interpreted as having the DNN in parallel with the BNN but doing a linear transformation of the DNN to the HF data before training the BNN. A more expressive and more flexible MF model is to avoid LR and just consider DNN-BNN, i.e. doing transfer-learning directly with a BNN on the HF data and disregarding the residual as in Eq. (3).
@@
In the DNN-LR-BNN model, the BNN models the residual as $r(\mathbf{x}^{h})=\mathbf{y}^{h}-\mathbf{m}(\mathbf{x}^{h})^{T}\bm{\rho}$, where the order of the basis $\mathbf{m}(\mathbf{x}^{h})$ is a hyperparameter that can be determined by cross-validation or hyperparameter optimization, similarly to the previously mentioned KRR-LR-GPR model. The most common prior assumed for BNNs is the Normal distribution and this choice also leads to the same distribution for the DNN-LR-BNN architecture because $r(\mathbf{x})$ converges to standard Normal distribution when the LF and HF models are highly correlated. There are many viable strategies for inference in BNNs, as reviewed previously: from Markov Chain Monte Carlo approaches like Hamiltonian Monte Carlo [52, 53] to Stochastic Gradient Langevin Dynamics (SGLD) [48], passing through Variational Inference approaches such as Bayes by Back-propagation [49], etc. In practice, for large datasets we recommend the use of a variant of SGLD called pSGLD [58] due to its superior scalability [59] – additional details are provided in Appendix A.1.1.
Contrary to the KRR-LR-GPR model, the determination of the transfer-learning parameters $\bm{\rho}$ is not done by Eq. (24) because it becomes intractable to calculate the covariance matrix of a BNN. Simultaneously, using cross-validation or other hyperparameter tuning strategies can be computationally intensive due to the time needed for BNN inference. Therefore, we propose to estimate $\bm{\rho}$ by solving the following optimization problem before training the BNN:
$\min_{\bm{\rho}}$ $\sum_{i=1}^{N^{h}}\left(\mathbf{y}^{h}-\mathbf{m(\mathbf{X}^{h})}^{T}\bm{\rho}\right)$ (8)
This equation is easy to optimize, akin to what is done in linear regression. Also, recall that by assuming $\rho_{0}=0$ and $\rho_{1}=1$, we obtain a vanilla MF model that fuses a DNN with a BNN without additional parameters. In summary, the main steps of DNN-LR-BNN are listed in Algorithm 2, leading to a Normal distribution prediction $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{s}_{f^{h}}^{2}(\mathbf{x}))$ at any unknown point.
@@
Result: $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{s}_{f^{h}}^{2}(\mathbf{x}))$
@@
   Step 2.1: Obtain $\bm{\rho}$ by minimizing $\sum_{i=1}^{n^{h}}\left(\mathbf{y}^{h}-\mathbf{m(\mathbf{X}^{h})}^{T}\bm{\rho}\right)$
@@ 4 Experiments
We analyze the performance of KRR-LR-GPR, DNN-LR-BNN and DNN-BNN models separately and compare them with state-of-the-art strategies. For comparison with the KRR-LR-GPR method, we select Co-Kriging [16], Hierarchical Kriging [23], and Scaled Kriging [26]. We consider 10 different functions to be learned using datasets with two fidelities, as described in Appendix C.1. These are low-dimensional functions with an input dimension spanning from $d=1$ to $d=8$, so they only need a small number of HF training samples. Importantly, the datasets utilized in this section are assumed to be noisy [^2] We implemented the existing methods such that noisy datasets could also be considered.. Additional experiments are conducted in Appendix D where noiseless datasets are also considered. Concerning the performance assessment of DNN-LR-BNN, comparison with other models is more challenging due to the multitude of hyperparameters and model choices that are possible when considering neural networks in different fidelities. Therefore, we compare DNN-LR-BNN with single-fidelity BNN and DNN-BNN models considering the same hyperparameters, as summarized in Appendix F. Several different numerical examples are considered.
Regarding the accuracy metrics, no single metric can objectively evaluate model performance for comparison. To this end, we considered Normalized Root Mean Square Error (NRMSE) and R2 Score to assess the global predictive mean performance, calculated based on the noiseless ground truth and predictive mean [60]. Concerning the uncertainty quantification metric, we adopted the Test Log-Likelihood (TLL) calculated based on the noisy test dataset [61]. In addition, the CPU execution time for LF and HF are recorded to evaluate computational efficiency, and each experiment is conducted on a node of an HPC cluster platform with an Intel® Xeon(R) E5-2643v3 CPU with 6 cores of 3.40GHz and 128 GB of RAM.
@@ 4.1.1 Illustrative example
Given the traditional importance of the Forrester function [16, 23], we start by illustrating the effectiveness of the KRR-LR-GPR method on this function – see Eq. (28) – and comparing it with alternative formulations seen in the literature that involve GPR models for both fidelities. We start by sampling 7 and 200 samples uniformly for HF and LF, respectively, and added Gaussian noise $\mathcal{N}(0,0.3^{2})$ on both LF and HF. However, our investigation later also considers different sample sizes and different noise levels. Fig.3 shows the results for all three MF-GPR models compared to the proposed KRR-LR-GPR model of an arbitrary realization.
@@
Table 2: Performance comparison between KRR-LR-GPR and other MF-GPR methods on the illustrative example
Methods NRMSE R2 Score TLL $\hat{\sigma}_{h}$ LF training time ($s$) HF training time ($s$)
Hierarchical Kriging 0.0912 0.9971 -0.5771 0.3610 4.6584 0.2015
Co-Kriging 0.0862 0.9974 -0.4991 0.3505 8.0879 0.3257
Scaled Kriging 0.2163 0.9842 -1.4960 1.1989 5.8713 0.2847
KRR-LR-GPR 0.1232 0.9949 -0.6676 0.4440 3.5073 0.1986
The Forrester function is easily predicted by all methods, except Scaled Kriging that exhibits lower performance. Note that the HF predictive means almost overlap perfectly with the HF truth with NRMSE approaching zero and R2 Score close to 1 in Table 2. Co-Kriging, Hierarchical Kriging, and KRR-LR-GPR all converge to similar $\hat{\rho}_{1}$, that is different from the imposed value of $\hat{\rho}_{1}=1$ for Scaled Kriging – affecting its performance. Notwithstanding, even this simple example already shows that the KRR-LR-GPR requires less training time than other MF-GPR approaches, due to the use of kernel ridge regression for the LF.
We then investigated the influence of imposing different correlations between the LF and HF data, as shown in Eq. (29), for the proposed KRR-LR-GPR method. Fig.4 presents the trained KRR-LR-GPR model considering different LF datasets. Three different LF functions are used to generate three different LF datasets, so the figures are labeled as “LF data 1”, “LF data 2” and “LF data 3” accordingly. Table 3 summarizes the performance of the KRR-LR-GPR model.
Figure 4: Predictions of KRR-LR-GPR for the Forrester function with different LF datasets of decreasing correlation (from left to right), as indicated in Table 3. The figure on the left corresponds to the single-fidelity result of GPR.
Table 3: Results comparison of the illustrative example with different LF datasets
@@
NRMSE 1.3129 0.0724 0.1226 0.2971
R2 Score 0.4188 0.9982 0.9949 0.9702
TLL -18513.24 -0.4253 -0.6637 -19161.03
hf training time ($s$) 0.1175 0.1635 0.1833 0.1080
lf training time ($s$) - 3.6946 4.3618 4.4272
$\hat{\sigma}^{h}$ 0.000064 0.3995 0.4411 0.000061
@@
Figure 5: Negative log marginal likelihood values of GPR and KRR-LR-GPR within parameter space. The contour plots only show regions where the negative log marginal likelihood value is smaller than 5 for better illustration, the maximum values for each method are more than $10^{7}$ The X-axis represents kernel parameter $\log_{10}(\theta)$ and the Y-axis is the estimated noise level $\hat{\sigma}^{h}$, the red point is the best parameter found for every method by "L-BFGS-B" with 10 restarts.
@@
From Fig.5 and Table 3 we see that the single-fidelity GPR cannot learn the underlying function because it fails to identify the noise level. However, the LF of the KRR-LR-GPR conditions the redidual of the GPR and leads to good predictions according to NRMSE and R2 Score evaluated. Unsurprisingly, performance deteriorates as the Pearson correlation coefficient $r$ between the two fidelities decreases from LF data 1 to 3. In terms of the estimated HF noise standard deviations, they are $0.3995$, $0.4411$, and $0.000061$, respectively. This can be understood by observing Fig.5 that shows the landscape of negative log marginal likelihood values in the parameter space for the three cases, where it is highlighted that in the third case there is no a clear optimum value (right most figure) and the marginal liklihood is more similar to the single fidelity GPR model (left most figure) – this explains why the noise level at the HF is not properly estimated. The results are intuitive and explainable: if the LF data has limited or nonexistent correlation with the HF, then the predictive capability of the MF model degrades.o
@@
Figure 6: Performance of KRR-LR-GPR on Forrester function with different LF samples and noise levels: the top and bottom rows have 11 and 200 LF samples, respectively, while all cases consider 7 HF samples. Noise level of LF indicated by $\sigma^{l}$ and number of LF samples is clear from the cross markers in each plot.
@@
Table 4: Results of KRR-LR-GPR on Forrester function with different LF samples and noise levels
LF samples $\sigma_{a}^{l}$ NRMSE R2 Score TLL $\hat{\sigma}^{h}$
11 0.0 0.0858 0.9975 -0.4811 0.2741
0.3 0.1971 0.9869 -0.9940 0.6089
0.5 0.3056 0.9685 -1.3586 0.9187
1.0 1.2941 0.4353 -47601.48 0.000061
200 0.0 0.0839 0.9976 -0.4715 0.2698
0.3 0.0815 0.9977 -0.4768 0.2428
0.5 0.0908 0.9972 -0.5523 0.2309
1.0 0.1355 0.9938 -0.7897 0.2597
@@
re, we also investigate the performance of the KRR-LR-GPR model with different noise levels (see Fig.6 and Table 4). We consider different Gaussian noise for the LF with standard deviations of $\sigma^{l}=0.0$, $\sigma^{l}=0.3$, $\sigma^{l}=0.5$, and $\sigma^{l}=1.0$, such that the robustness of KRR-LR-GPR to LF noise is reported while keeping the HF noise as $\sigma^{h}=0.3$. The figure and table clarify that the model predicts both the function and noise level accurately for the cases with enough LF data. Conversely, it struggles to identify proper noise levels in cases of 11 LF points and higher LF noise levels, as expected.
4.1.2 Comprehensive experiments of KRR-LR-GPR
Figure 7: Comparisong of KRR-LR-GPR with state-of-the-art MF methods for different sizes of HF datasets (here showing results for the Booth function, $d=2$, Pearson cor. coef. $r=0.925$ ). Different colors represent different methods. To interpret the results, note that a smaller NRMSE or a higher TLL is better. Note the significantly faster LF training time for KRR-LR-GPR compared to others.

### v2

@@ 3.1 Model for data-scarce or low-dimensional scenarios: KRR-LR-GPR
We argue that a logical LF model to pair with GPRs in a data-scarce HF scenario is KRR, i.e. the deterministic formulation of GPR [16]. KRR is a kernel machine learning method that is a point estimate of a GPR with the same kernel, therefore also equivalent to a DNN with an infinitely wide hidden layer [55]. KRR is robust to noisy data, and easy to train due to having few hyperparameters. More importantly, existing KRR approaches can lower complexity to a range between $\mathcal{O}(N^{2})$ and $\mathcal{O}(N)$, thus handling larger datasets than PCE or GPR [56, 57] – invaluable for LF datasets because they are typically larger than HF ones. Other LF models such as LR or DNNs (e.g., see Table 1) have important drawbacks. LR models are sensitive to the choice of the basis functions (often considered to be polynomials whose order is a hyperparameter); this is why we argue for their use as a simple transfer-learning model $g(\mathbf{x})$ but not as a LF or residual model. Conversely, DNNs have a large number of hyperparameters and are therefore less practical to train. Good practice involves starting with simpler models such as KRR that are easier to train and only considering DNNs when the simpler methods fail. We note that GPRs can also be a good choice for a LF model if the LF dataset is small, but training and inference time increases when compared to KRR, as shown later.
@@
The proposed KRR-LR-GPR model established on Equation 4 is easy to implement and the main steps are listed in Algorithm 1, while additional details are provided in Appendix B. Recall that the key idea is to train a GPR model where, instead of considering a zero mean function, we assume it to be a linear model whose features are the LF model. This is advantageous as there is an explicit solution for finding the parameters $\boldsymbol{\rho}$ from the HF data, no longer treating them as hyperparameters: Step 2.1 in Algorithm 1 or Equation 22 in Appendix B. The mean model of the GPR then captures most of the HF response by linear transfer-learning of the LF model, leaving the remaining nonparametric approximation for the residual via a Gaussian process whose kernel hyperparameters are optimized as usual (Step 2.2 in Algorithm 1). The resulting KRR-LR-GPR model predicts the response and corresponding uncertainty, usually by following a Normal distribution $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{\sigma}_{h}^{2}(\mathbf{x}))$ to ensure fast training and inference (exact integration of posterior and posterior predictive distributions). We find that a first-order linear transfer-learning model including a bias term, i.e. $\mathbf{m}(\mathbf{x})=[1,f^{l}(\mathbf{x})]^{T}$, provides robust results according to the ablation study for different problems (see Section E.1).
@@
Result: $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{\sigma}_{h}^{2}(\mathbf{x}))$
@@ 3.2 Model for data-rich or high-dimensional scenarios: DNN-LR-BNN
Pairing a DNN as the LF model with a BNN in a MF model leverages the scalability of DNNs with the predictive abilities of BNNs. We propose a novel configuration summarized in Figure 2 labeled DNN-LR-BNN, where transfer-learning is done via LR as in Equation 4, and the residual is modeled by a BNN.
@@
In the DNN-LR-BNN model, the BNN learns the residual $r(\mathbf{x}^{h})=\mathbf{y}^{h}-\mathbf{m}(\mathbf{x}^{h})^{T}\bm{\rho}$, where the order of $\mathbf{m}(\mathbf{x}^{h})$ is a hyperparameter. Usually, we assume Normal distributions for the prior and observation distribution of BNNs, leading to a Normal PPD; Therefore, the DNN-LR-BNN model also has a Normal PPD. As mentioned previously, there are many viable strategies for inference in BNNs: from Markov Chain Monte Carlo approaches like Hamiltonian Monte Carlo [51, 52] to SGLD [47], passing through Variational Inference approaches such as Bayes by Back-propagation [48]. In practice, for large datasets, we recommend the use of a variant of SGLD called pSGLD [58] due to its superior scalability [59] – additional details are provided in Section A.1.1.
Contrary to the KRR-LR-GPR model, the determination of the transfer-learning parameters $\bm{\rho}$ is not done by Equation 22 because it becomes intractable to calculate the covariance matrix of a BNN. Simultaneously, using cross-validation or other hyperparameter tuning strategies can be computationally intensive due to the time needed for BNN inference. Therefore, we propose to estimate $\bm{\rho}$ by solving the following optimization problem before training the BNN:
$\hat{\bm{\rho}}=\arg\min_{\bm{\rho}}\sum_{i=1}^{N^{h}}\left\|\mathbf{y}^{h}-\mathbf{m}(\mathbf{X}^{h})^{T}\bm{\rho}\right\|^{2}$ (6)
This equation is easy to optimize, akin to what is done in linear regression. Note that if we assumed $\rho_{0}=0$ and $\rho_{1}=1$, a vanilla MF model would be obtained (fusing a DNN with a BNN without additional parameters). The main steps of DNN-LR-BNN are listed in Algorithm 2, leading to a prediction following a Normal distribution $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{\sigma}_{h}^{2}(\mathbf{x}))$ at any unknown point.
@@
Result: $\mathcal{N}(\hat{f}^{h}(\mathbf{x}),\hat{\sigma}_{h}^{2}(\mathbf{x}))$
@@
   Step 2.1: Obtain $\bm{\rho}$ by minimizing Equation 6
@@ 4 Experiments
We separately analyze the performance of KRR-LR-GPR and DNN-LR-BNN models and compare them with state-of-the-art strategies. For comparison with the KRR-LR-GPR method, we select Co-Kriging [16], Hierarchical Kriging [22], and Scaled Kriging [26]. We first consider 10 different functions to be learned using datasets with two fidelity levels, as described in Section C.1. These are low-dimensional functions with an input dimension spanning from $d=1$ to $d=8$, so they only need a small number of HF training samples. Importantly, the datasets utilized in this section are assumed to be noisy [^2] We implemented the existing methods such that noisy datasets could also be considered.. Additional experiments are conducted in Appendix D where noiseless datasets are also considered. Then, we train KRR-LR-GPR to predict the aerodynamic coefficients for a NACA 0012 airfoil in Section 4.1.3. Concerning the performance assessment of DNN-LR-BNN, comparison with other models is more challenging due to the multitude of hyperparameters and model choices that are possible when considering neural networks in different fidelity levels. Therefore, we compare DNN-LR-BNN with single-fidelity BNN and DNN-BNN models [1] considering the same hyperparameters, as summarized in Appendix F. Several different numerical examples are considered in Section 4.2.1 and Section 4.2.2; and we finally validate its performance with a material structure-property linkages problem in Section 4.2.3.
For model comparison, considering that no single metric can objectively evaluate model performance, we use NRMSE and R2 Score to assess the predicted mean performance [60]. Concerning the uncertainty quantification metric, we adopt the TLL when we only have access to the noisy test dataset, as occurs in practice [61]. In addition, we implemented all the algorithms by ourselves and in the same coding language to record the CPU execution time for LF and HF. Each experiment is conducted on a node of an HPC cluster platform with an Intel® Xeon(R) E5-2643v3 CPU with 6 cores of 3.40GHz and 128 GB of RAM. Rigorous comparison of computational cost is challenging, as each method can be optimized differently. Our implementations and results are made publicly available, in an attempt to demonstrate fairness in our comparisons and motivate future research on this topic.
@@ 4.1.1 Illustrative example
Given the traditional importance of the Forrester function [16, 22], we start by illustrating the effectiveness of the KRR-LR-GPR method for this example – see Equation 26. First, we compare the proposed method with methods in the literature that involve GPR models for both fidelity levels. We begin in Figure 3 and Figure 4 by considering LF and HF datasets with 200 and 7 uniformly sampled points, respectively. Note that our datasets include Gaussian noise $\mathcal{N}(0,0.3^{2})$ on both fidelity levels (LF and HF). Figure 6 considers different LF dataset sizes and different noise levels. Unsurprisingly, due to the simplicity of the Forrester function and its low-dimensionality, this first example was found to be trivial for most methods.
@@
Table 2: Performance comparison between KRR-LR-GPR and other MF-GPR methods on the illustrative example over 5 independent runs.
Methods NRMSE R2 Score TLL $\hat{\sigma}_{h}$ Training time($s$)
LF HF
Hierarchical Kriging 0.1408 (±$\pm 0.044$) 0.9927 (±$\pm 0.005$) -0.7895 (±$\pm 0.189$) 0.4360 (±$\pm 0.167$) 4.2705 (±$\pm 1.048$) 0.2429 (±$\pm 0.056$)
Co-Kriging 0.1429 (±$\pm 0.041$) 0.9925 (±$\pm 0.004$) -0.9651 (±$\pm 0.468$) 0.3446 (±$\pm 0.178$) 4.2305 (±$\pm 0.676$) 0.4458 (±$\pm 0.078$)
Scaled Kriging 0.2435 (±$\pm 0.032$) 0.9796 (±$\pm 0.005$) -1.3306 (±$\pm 0.359$) 0.5284 (±$\pm 0.485$) 4.0597 (±$\pm 0.792$) 0.2202 (±$\pm 0.038$)
KRR-LR-GPR 0.1455 (±$\pm 0.047$) 0.9922 (±$\pm 0.005$) -0.8573 (±$\pm 0.308$) 0.3803 (±$\pm 0.164$) 2.4128 (±$\pm 1.348$) 0.2422 (±$\pm 0.037$)
Figure 3 shows the results for all three MF-GPR models compared to the proposed KRR-LR-GPR model (for one of the 5 realizations chosen at random). Table 2 summarizes the performance metrics obtained from 5 independent runs with random seeds for each method. As mentioned, the Forrester function is easily predicted by all methods, except Scaled Kriging. The HF predicted means overlap almost perfectly with the HF ground-truth, as quantified in Table 2 by an NRMSE approaching zero and an R2 Score close to 1. Co-Kriging, Hierarchical Kriging, and KRR-LR-GPR all converge to similar $\hat{\rho}_{1}$. Note that Scaled Kriging considers $\hat{\rho}_{1}=1$, leading to inferior predictions. Unsurprisingly, KRR-LR-GPR requires less training time than all MF-GPR methods due to the use of KRR for the LF.
Figure 4 and Table 3 summarize the performance of KRR-LR-GPR when considering datasets with different correlations between LF and HF data. Three different LF functions are used to generate three different LF datasets, as explained in Section C.1, leading to the labels LF data 1, LF data 2 and LF data 3 in Figure 4.
Figure 4: Predictions of KRR-LR-GPR for the Forrester function with different LF datasets of decreasing correlation with the HF dataset (from left to right), as indicated in Table 3. Different correlations between LF and HF data are obtained according to Equation 27, where the parameters $A$, $B$, and $C$ of the LF function are changed as follows: LF data 1 considers $A=1,B=0,C=5$; LF data 2 considers $A=0.5,B=10,C=5$; LF data 3 considers $A=0.1,B=10,C=0.1$. The figure on the left corresponds to the single-fidelity result of GPR.
Table 3: Summary of performance of KRR-LR-GPR model considering different LF datasets over 5 independent runs.
@@
NRMSE 1.3176 (±$\pm 0.006$) 0.0823 (±$\pm 0.024$) 0.1216 (±$\pm 0.025$) 0.6524 (±$\pm 0.057$)
R2 Score 0.4147 (±$\pm 0.005$) 0.9975 (±$\pm 0.001$) 0.9948 (±$\pm 0.002$) 0.8556 (±$\pm 0.024$)
TLL -3.7502 (±$\pm 1.490$) -0.5303 (±$\pm 0.300$) -0.8650 (±$\pm 0.391$) -2.2474 (±$\pm 0.094$)
Training time ($s$) HF 0.1560 (±$\pm 0.031$) 0.1731(±$\pm 0.023$) 0.1707 (±$\pm 0.016$) 0.1398 (±$\pm 0.039$)
LF 1.5543 (±$\pm 0.562$) 1.3730 (±$\pm 0.326$) 1.2043 (±$\pm 0.523$)
$\hat{\sigma}_{h}$ 0.000060 (±$\pm 0.000001$) 0.2851 (±$\pm 0.050$) 0.2447 (±$\pm 0.082$) 2.6200 (±$\pm 0.307$)
@@
Figure 5: Negative log marginal likelihood values of GPR and KRR-LR-GPR within parameter space. The contour plots only show regions where the negative log marginal likelihood value is smaller than 5 for better illustration, the maximum values for each method are more than $10^{7}$. The X-axis represents kernel parameter $\log_{10}(\theta)$ and the Y-axis is the estimated noise level $\hat{\sigma}_{h}$, the red point is the best parameter found for every method by "L-BFGS-B" with 10 restarts.
@@
Figure 4 and Table 3 clarify that the single-fidelity GPR model cannot learn the underlying function well because it fails to identify the noise level. However, the transfer-learned LF model of KRR-LR-GPR provides appropriate conditioning of the GPR residual and leads to good predictions (see appropriate performance scores, NRMSE, and R2 Score). Performance deteriorates as the Pearson correlation coefficient $r$ between the two fidelity levels decreases from LF data 1 to 3. Similarly, model uncertainty at the HF is lower when the LF datasets are more correlated with the HF dataset (higher Pearson correlation coefficient). Figure 5 also explains this by showing the negative log marginal likelihood landscape in the parameter space for a particular run of each of the three cases, where it is highlighted that in the third case (LF data 3) there is no clear optimum value (rightmost figure), exhibiting a similar marginal likelihood to the single-fidelity GPR model (leftmost figure) – this explains why the noise level at the HF is not properly estimated. The results are intuitive: if the LF data has limited or nonexistent correlation with the HF, then there is a degradation of the prediction quality of the MF model.
@@
Figure 6: Performance of KRR-LR-GPR model on Forrester function with different number of LF samples and noise levels: the top and bottom rows show results obtained for a LF dataset with 11 and 200 samples, respectively. In this figure, the HF dataset is not changed and considers 7 samples. The noise level of LF is indicated by $\sigma_{l}$. The number of LF samples is clear from the cross markers in each plot.
@@
Table 4: Results of KRR-LR-GPR on Forrester function with different LF samples and LF noise levels over 5 independent runs
LF samples $\sigma_{l}$ NRMSE R2 Score TLL $\hat{\sigma}_{h}$
11 0.0 0.1267 (±$\pm 0.012$) 0.9945 (±$\pm 0.001$) -0.6279 (±$\pm 0.007$) 0.3300 (±$\pm 0.052$)
0.3 0.2572 (±$\pm 0.088$) 0.9751 (±$\pm 0.018$) -11304.37 (±$\pm 14787.37$) 0.3229 (±$\pm 0.378$)
0.5 0.3184 (±$\pm 0.073$) 0.9640 (±$\pm 0.016$) -11202.4 (±$\pm 14779.75$) 0.4992 (±$\pm 0.523$)
1.0 0.4125 (±$\pm 0.139$) 0.9361 (±$\pm 0.045$) -11128.88 (±$\pm 14842.75$) 0.6761 (±$\pm 0.682$)
200 0.0 0.1184 (±$\pm 0.015$) 0.9952 (±$\pm 0.001$) -0.6106 (±$\pm 0.092$) 0.3374 (±$\pm 0.057$)
0.3 0.1119 (±$\pm 0.016$) 0.9957 (±$\pm 0.001$) -0.5398 (±$\pm 0.080$) 0.3159 (±$\pm 0.017$)
0.5 0.1184 (±$\pm 0.020$) 0.9951 (±$\pm 0.001$) -0.5693 (±$\pm 0.094$) 0.3136 (±$\pm 0.061$)
1.0 0.1563 (±$\pm 0.026$) 0.9915 (±$\pm 0.003$) -0.7243 (±$\pm 0.121$) 0.3731 (±$\pm 0.055$)
@@
Furthermore, we investigate the performance of the KRR-LR-GPR model with different noise levels and different LF dataset sizes (see Figure 6 and Table 4). We consider different Gaussian noise for the LF with standard deviations of $\sigma_{l}=0.0$, $\sigma_{l}=0.3$, $\sigma_{l}=0.5$, and $\sigma_{l}=1.0$, such that the robustness of KRR-LR-GPR to LF noise is assessed while maintaining the HF noise as $\sigma_{h}=0.3$. The figure and table clarify that the model predicts both the function and noise level accurately for the cases with enough LF data. Conversely, it struggles to identify proper noise levels when the LF dataset only has 11 points and for higher LF noise levels, as expected.
4.1.2 Comprehensive numerical experiments of KRR-LR-GPR